The quality control manager of WHEAT Bakery has observed (over a long period of time) that 2% of loaves of bread baked in the evening are either underweight or over weight and they must not be sold. If she took a sample of 6 such loaves of bread for inspection, find the probability that:


Exactly 2 are unacceptable.

At most 3 are unacceptable.

Between 3 and 5 (inclusive) are no unacceptable.


(b) How many loaves of bread would be expected to be unacceptable from a sample of 100 such loaves?

(c) Suppose that A and B are events in a sample space S such that p ( A )  0 .4 , p(A  B)  0.7 and p(B)  k

(i) What is the numerical value of k for which events A and B are independent.

(ii) What is the numerical value of k for which A and B are mutually exclusive.


1
Expert's answer
2021-11-24T17:51:56-0500

(a) Let "X=" the number of unacceptable loaves of bread: "X\sim Bin(n, p)."

Given "n=6, p=0.02,q=1-0.02=0.98"

(i) Exactly 2 are unacceptable.


"P(X=2)=\dbinom{6}{2}(0.02)^2(0.98)^{6-2}"

"=0.00553420896"

(ii) At most 3 are unacceptable.


"P(X\leq 3)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)=\dbinom{6}{0}(0.02)^0(0.98)^{6-0}"

"+\dbinom{6}{1}(0.02)^1(0.98)^{6-1}+\dbinom{6}{2}(0.02)^2(0.98)^{6-2}"

"+\dbinom{6}{3}(0.02)^3(0.98)^{6-3}=0.885842380864"


"+0.108470495616+0.00553420896"

"+0.00015059072=0.99999767616"



(iii) Between 3 and 5 (inclusive) are no unacceptable. That is at most 2 are unacceptable or Exactly 6 are unacceptable.


"P(X\leq 2\ or\ X=6)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=6)"

"=\dbinom{6}{0}(0.02)^0(0.98)^{6-0}+\dbinom{6}{1}(0.02)^1(0.98)^{6-1}"

"+\dbinom{6}{2}(0.02)^2(0.98)^{6-2}+\dbinom{6}{6}(0.02)^6(0.98)^{6-6}"

"=0.885842380864+0.108470495616"

"+0.00553420896+0.000000000064"

"=0.999847085504"

(b)

Given "n=100, p=0.02"


"E(X)=np=100(0.02)=2"

2 loaves of bread would be expected to be unacceptable from a sample of 100 such loaves.


(c)

Given "P(A)=0.4, P(A\cup B)=0.7, P(B)=k"



"P(A\cup B)=P(A)+P(B)-P(A\cap B)"


"P(A\cap B)=P(A)+P(B)-P(A\cup B)"

"P(A\cap B)=0.4+k-0.7=k-0.3"

(i)


"P(A\cap B)=P(A)P(B)"

"k-0.3=0.4k"

"0.6k=0.3"

"k=0.5"

(ii)


"P(A\cup B)=P(A)+P(B)"

"0.7=0.3+k"

"k=0.4"